OpenStudy (anonymous):

Find the absolute max and min as well as all the values of x where they occur using the given function and domain. X=X+e^-5X

OpenStudy (anonymous):

i assume the left hand side is \(f(x)\)

OpenStudy (anonymous):

The derivative I got from this is (1+e^-5x) (-5)

OpenStudy (anonymous):

as in \[f(x)=x+e^{-5x}\]

OpenStudy (anonymous):

Yes, sorry.

OpenStudy (anonymous):

your derivative is wrong i think

OpenStudy (anonymous):

should just be \[f'(x)=1-5e^{-5x}\]

OpenStudy (anonymous):

How do I solve that to find the critical points? How do I make the e go away?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

set \[1-5e^{-5x}=0\]

OpenStudy (anonymous):

then go to \[5e^{-5x}=1\] divide by \(5\) and get \[e^{-5x}=\frac{1}{5}\]

OpenStudy (anonymous):

now get rid of the \(e\) by rewriting in logarithmic form as \[-5x=\ln(\frac{1}{5})\]

OpenStudy (anonymous):

or if you prefer \[-5x=\ln(.2)\] then divide and get \[x=-\frac{\ln(.2)}{5}\]

OpenStudy (anonymous):

that is how you make the \(e\) go away, by taking the log

OpenStudy (anonymous):

Thank you! You've been such a big help. I always thought though that on equations that had e in it, like the one we just did, that you still took the derivative of the x. Lol.

OpenStudy (anonymous):

you do if you need to you have to use the chain rule, so for example if \[f(x)=e^{x^2}\] then \[f'(x)=2xe^{x^2}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

But not when it's connected to something like 1+e to the whatever?

OpenStudy (anonymous):

i am not exactly sure what you mean...

OpenStudy (anonymous):

I figured it out. Lol